# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 3. $(3)$
    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{(n)}\|_{\mathbb{B}}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1,$
    and for any $j\geq k$, there exists a constant $\tilde{c}>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|{\tilde{v}}_{t}^{(n)}|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+\tilde{c}{\|\phi\|_{L^{k}(\mu)}^{j}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k},n\geq 1.$
Indeed, by (i), the fixed points are consistent:
$\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{t\land\tau_{n}}=\big{(}v^{(n+i)},{\tilde{v}}^{(n+i)}\big{)}_{t\land\tau_{n}},\ \ n,i\geq 1,t\in[0,T],$
so that $\mathbb{P}$-a.s.
$\big{(}v(\mu,\phi),\tilde{v}(\mu,\phi)\big{)}:=\lim_{n\rightarrow\infty}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}=\sum_{n=1}^{\infty}1_{[\tau_{n-1},\tau_{n})}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}_{\cdot\land\tau_{n}}$
exists in $C([0,T];\mathbb{B}\times\mathbb{R}^{d})$, and by (ii) the estimates (3.16) and (3.17) hold. Combining this with
(3.21) $\Phi^{(n)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}:=\big{(}\bar{H},H_{\cdot\land\tau_{n}}\big{)}\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)}=\big{(}v^{(n)},{\tilde{v}}^{(n)}\big{)},\ \ n\geq 1,$
we see that $(v(\mu,\phi),\tilde{v}(\mu,\phi))$ solves (3.14), and the uniqueness follows from that of the fixed point for $\Phi^{(n)},n\geq 1.$ Next, we will prove (i) and (ii) one by one.
@@
(3.22) $\bar{H}:\mathscr{V}_{0}\times\mathscr{C}_{k}\rightarrow\mathscr{V}_{0},$
so that $\Phi^{(n)}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$ for every $n\geq 1$.
By (3.3), we have
(3.23) $I_{t}^{V}(\mu,\phi)=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}\big{\{}V(X_{t}^{x,\mu})\big{\}}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).$
Similar to (2.11)-(2.13), for any $\lambda\geq\lambda_{0}$, $\gamma\in\mathscr{P}_{k}$ and $x\in\mathbb{R}^{d}$, consider
$\text{\rm{d}}\theta^{\lambda,\gamma,B}_{t}(x)=B_{t}((\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\theta^{\lambda,\gamma,B}_{0}(x)=\Theta^{\lambda,\gamma}_{0}(x),$
and let
$\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)=(\Theta^{\lambda,\gamma}_{t})^{-1}(\theta^{\lambda,\gamma,B}_{t}(x)),\ \ t\in[0,T],x\in\mathbb{R}^{d}.$
Then we have
$\text{\rm{d}}\Theta^{\lambda,\gamma}_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x))=B_{t}(\tilde{\theta}^{\lambda,\gamma,B}_{t}(x),P_{t}^{\ast}\gamma,P_{t}V(\gamma))\text{\rm{d}}t,\ \ t\in[0,T],\tilde{\theta}^{\lambda,\gamma,B}_{0}(x)=x\in\mathbb{R}^{d}.$
Repeating the proof of Lemma 2.3, we conclude that for any $p\geq 1$, there exists a constant $c_{p}>0$ such that
(3.24) $\mathbb{E}[|X^{x,\gamma}_{t}-\tilde{\theta}^{\lambda,\gamma,B}_{t}(x)|^{p}]\leq c_{p}t^{\frac{p}{2}},\ \ t\in[0,T],x\in\mathbb{R}^{d},\gamma\in\mathscr{P}_{k},$
@@
By $[V]_{\alpha}\leq 1$, (2.22) in Lemma 2.5 for $p=2$ and $y=\tilde{\theta}^{\lambda_{0},\mu,B}_{t}(x)$, (3.25), (2.1), (3.24) and (3.23), we find a constant $c_{1}>0$ such that
(3.26) $\begin{split}\|I_{t}^{V}(\mu,\phi)\|_{\mathbb{B}}&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}}.\end{split}$
To estimate $\tilde{I}_{t}^{V}(\mu,v,\tilde{v})$, noting that $X_{t}^{\mu}$ solves (3.9) for $\gamma=\mu$ and initial value $X_{0}^{\mu}$ replacing $x$, by the standard Markov property for solutions of (3.9), we have
$\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}V(X_{t}^{x,\mu})\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}}$

### v4

@@ 2. $(2)$
    $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}^{\mu,\phi}\|_{\infty}\leq c{\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ (t,\mu)\in[0,T]\times\mathscr{P}_{k}.$
Next, we will prove (i) and (ii) one by one.
@@
(3.20) $H:\mathscr{V}_{0}\rightarrow\mathscr{V}_{0}.$
Recall that $\theta^{\lambda,\gamma}$ and $\tilde{\theta}^{\lambda,\gamma}$ are defined in (2.15)-(2.17). Since (B) implies (A), we conclude that (2.18) still holds.
By $[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)]_{\alpha}\leq 1$ due to $(B_{2})$, (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x)$, (3.3), (2.18), (2.1) and (3.15), we find a constant $c_{1}>0$ such that
(3.21) $\begin{split}|I_{t}^{\mu,\phi}(y,\nu)|&\leq\frac{c_{1}}{\sqrt{t}}\int_{\mathbb{R}^{d}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}|\phi(x)|\mu(\text{\rm{d}}x)}\\ &{\leq\frac{c_{1}}{\sqrt{t}}\alpha\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)},\ \ t\in(0,T],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}},y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
So, by (3.14), $(B_{2})$, (2.10) in Lemma 2.3 for $p=2$ and $z=\tilde{\theta}^{\lambda,\mu}_{t}(x)$, (2.1) and (2.18), we find a constant $c_{2}>0$ such that
(3.22) $\begin{split}&|\tilde{I}_{t}^{V}(y,\nu)|\leq c_{2}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}{\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}}\|V_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
Combining this with (3.19) and (3.21), we find a constant $c_{3}>0$ such that
(3.23) $\begin{split}\|\{H(V)\}_{t}\|_{\infty}\leq&\,c_{3}\|\phi\|_{L^{k}(\mu)}\sqrt{\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}+c_{3}\sqrt{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}\big{(}s^{\frac{1}{2}}\big{)}}\|V_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}.\end{split}$
Then (3.20) follows from the fact that (2.40) implies
(3.24) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0.$
(b) We intend to prove that $H$ in (i) has a unique fixed point in $\mathscr{V}_{0}.$ Obviously, for any $\delta>0$, $\mathscr{V}_{0}$ is complete under the metric
$\rho_{\delta}\big{(}V,U\big{)}:=\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta t}\|V_{t}-U_{t}\|_{\infty},\ \ V,U\in\mathscr{V}_{0}.$
So, it suffices to prove the contraction of $H$ in $\rho_{\delta}$ for large enough $\delta>0$.
By (3.19), (3.21) and (3.22), we find a constant $c_{4}>0$ such that
|$|\{H(V)\}_{t}(y,\nu)-\{H(U)\}_{t}(y,\nu)|=\frac{\{t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\}^{\frac{1}{2}}}{\alpha\big{(}t^{\frac{1}{2}}\big{)}}|\tilde{I}_{t}^{V-U}(y,\nu)|$
$\leq c_{4}\bigg{(}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|V_{s}-U_{s}\|^{2}_{\infty}\text{\rm{d}}s\bigg{)}^{\frac{1}{2}},\ \ V,U\in\mathscr{V}_{0},t\in[0,T].$
Combining this with (3.24), we conclude that $H$ is contractive in the complete metric space $(\mathscr{V}_{0},\rho_{\delta})$ for large enough $\delta>0$, and hence has a unique fixed point denoted by $v^{\mu,\phi}$.
(2) Proof of (ii). By (3.23) and noting that $H(v^{\mu,\phi})=v^{\mu,\phi}$, we derive
$\|v_{t}^{\mu,\phi}\|_{\infty}^{2}\leq 2c_{3}^{2}\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\|\phi\|_{L^{k}(\mu)}^{2}+2c_{3}^{2}t\tilde{\alpha}\big{(}t^{\frac{1}{2}}\big{)}\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(s^{\frac{1}{2}})}\|v_{s}^{\mu,\phi}\|^{2}_{\infty}\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (3.24) and Gronwall’s inequality, we find a constant $c_{5}>0$ such that for any $t\in[0,T]$,
$\|v_{t}^{\mu,\phi}\|_{\infty}\leq c_{5}\|\phi\|_{L^{k}(\mu)}\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})},\ \ \mu\in\mathscr{P}_{k},\ \phi\in T_{\mu,k}.$
This proves (ii). ∎
3.3 Proof of Theorem 3.3
By Lemma 3.4, the proof of Theorem 3.3(1) is completed by the following lemma.
Lemma 3.5.
Assume (B). Then for any $\mu\in\mathscr{P}_{k}$, $\phi\in T_{\mu,k}$, the function $h:(0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}\rightarrow\mathbb{R}^{d}$ defined by
$h_{t}(y,\nu):=D_{\phi}^{I}\big{\{}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}},\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$
[…]
$\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\bigg{\{}\sup_{t\in(0,T]}\frac{\sqrt{t}}{\alpha(t^{\frac{1}{2}})}\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|D_{\phi}^{I}\big{\{}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}}|\bigg{\}}\leq c.$
Proof.
(a) By Lemma 3.4, it suffices to prove (3.17). For simplicity, for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, let
(3.25) $U_{t}(y,\nu,z):=D^{E}b_{t}^{(1)}(y,\nu)(z),\ \ z\in\mathbb{R}^{d}.$
Moreover, simply denote
$v_{t}^{\varepsilon}(y,\nu):=v_{t}^{\varepsilon,1}(y,\nu)+v_{t}^{\varepsilon,2}(y,\nu),$
$v_{t}^{\varepsilon,1}(y,\nu):=\frac{P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)}{\varepsilon},$
$v_{t}^{\varepsilon,2}(y,\nu):=\frac{P_{t}^{\mu_{\varepsilon}}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})}{\varepsilon}.$
Next, for $v_{t}^{\mu,\phi}$ in Lemma 3.4, let
(3.26) $\begin{split}&\hat{v}_{t}(y,\nu):={\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}v_{t}^{\mu,\phi}(y,\nu),\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k},\end{split}$
@@
(3.27) $\hat{v}_{t}^{1}(y,\nu):=I_{t}^{U_{t}(y,\nu,\cdot)}(\mu,\phi),\ \ \hat{v}_{t}^{2}(y,\nu):=\hat{v}_{t}(y,\nu)-I_{t}^{U_{t}(y,\nu,\cdot)}(\mu,\phi).$
Noting that
$P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu_{\varepsilon})-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(\mu)$
(3.28) $=\int_{\mathbb{R}^{d}}\Big{[}P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x+\varepsilon\phi(x))-P_{t}^{\mu}U_{t}(y,\nu,\cdot)(x)\Big{]}\mu(\text{\rm{d}}x),$
by (3.4) for $f=U_{t}(y,\nu,\cdot)$, we obtain
(3.29) $\lim_{\varepsilon\rightarrow 0}\big{|}v_{t}^{\varepsilon,1}(y,\nu)-\hat{v}_{t}^{1}(y,\nu)\big{|}=0,\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.$
Since (3.16) holds for $V_{t}=v_{t}^{\mu,\phi}$, (3.25)-(3.27) imply that